OpenStudy (anonymous):

evaluate the indefinite integral. sec(x)^21tan(x)dx

OpenStudy (anonymous):

let u=tan(x)

OpenStudy (anonymous):

du= sec^2 x dx

OpenStudy (anonymous):

= integral udu = (1/2)u^2 +C = (1/2)tan^2(x) +C